OpenStudy (anonymous):

for f(x)=x+10 (fof^-1)(10)

OpenStudy (anonymous):

It's been a long time since I have worked with them; but here's my solution: f(x) = x + 10 => x = f(x) - 10 => f(x)^-1 = x - 10 fof^-1(x) = x - 10 + 10 = x fof^-1 (10) = 10